# ORACLE-Grasp: Zero-Shot Task-Oriented Robotic Grasping using Large Multimodal Models

### v1

@@ Abstract
Grasping unknown objects in unstructured environments remains a fundamental challenge in robotics, requiring both semantic understanding and spatial reasoning. Existing methods often rely on dense training datasets or explicit geometric modeling, limiting their scalability to real-world tasks. Recent advances in Large Multimodal Models (LMMs) offer new possibilities for integrating vision and language understanding, but their application to autonomous robotic grasping remains largely unexplored. We present ORACLE-Grasp, a zero-shot framework that leverages LMMs as semantic oracles to guide grasp selection without requiring additional training or human input. The system formulates grasp prediction as a structured, iterative decision process, using dual-prompt tool calling to first extract high-level object context and then select task-relevant grasp regions. By discretizing the image space and reasoning over candidate areas, ORACLE-Grasp mitigates the spatial imprecision common in LMMs and produces human-like, task-driven grasp suggestions. Early stopping and depth-based refinement steps further enhance efficiency and physical grasp reliability. Experiments demonstrate that the predicted grasps achieve low positional and orientation errors relative to human-annotated ground truth and lead to high success rates in real-world pick up tasks. These results highlight the potential of combining language-driven reasoning with lightweight vision techniques to enable robust, autonomous grasping without task-specific datasets or retraining.
I Introduction
Robotic grasping remains a core challenge in autonomous manipulation, especially in unstructured and novel environments. Traditional approaches often depend on precise object models, depth-based geometry recognition, or supervised learning pipelines trained on large-scale datasets [1, 2]. However, real-world deployment frequently requires interacting with novel objects exhibiting unpredictable shapes with task-based reasoning. In such scenarios, accurate grasp localization, identifying grasp points and orientations that are both feasible and task-appropriate, becomes critical [3]. While recent deep learning methods have advanced grasp quality prediction [4], they primarily focus on geometric properties and neglect semantic reasoning about the object or intended task. This gap motivates the development of new frameworks that combine high-level semantic understanding with spatial reasoning, enabling more human-like, adaptive grasp selection in the real-world.
@@ I Introduction
Previous approaches have typically relied on geometric analysis of depth data [5, 6], model-based grasp synthesis [7], or extensive training of deep networks on large-scale grasp datasets [8, 9]. Early work on learning-based methods for grasping novel objects, used machine learning techniques to predict effective grasp points directly from 2D images without requiring complete 3D reconstruction or object recognition [10]. However, deep learning-based approaches typically focus on learning and predicting the quality of candidate grasps, without integrating semantic reasoning about the task or the object’s context [11]. While recent methods have made significant advances in pixel-wise grasp prediction and affordance learning [12, 13], they often lack the semantic reasoning needed for task-oriented grasping, especially in zero-shot settings where the object has not been seen before.
Large Language Models (LLMs) have significantly advanced robotics by enabling natural language understanding for task planning, human-robot interaction, and high-level decision-making [14, 15]. These models allow robots to interpret abstract instructions (e.g., ”organize the desk”) and generate structured action sequences with explainable reasoning. Building on these capabilities, recent developments have introduced Large Multimodal Models (LMMs), which integrate vision and language processing to tackle embodied tasks [16, 17]. By jointly reasoning over visual inputs and textual prompts, LMMs extend the functionality of LLMs, enabling robots to perceive, interpret, and act within their environments with greater semantic understanding.
Prominent examples of LMMs include CLIP [18], Flamingo [19], and LLaVA [20], which have demonstrated impressive capabilities in grounding visual scenes through natural language reasoning [21]. These models are typically trained on large-scale datasets comprising paired image and text data, allowing them to align visual and linguistic representations within a shared semantic space. As a result, LMMs can generalize across domains, leverage contextual information, and infer high-level semantic relationships that are critical for flexible and task-driven robotic behaviors [22]. Recently, LMMs have begun to be applied in robotics to bridge perception, reasoning, and action [16, 23, 24], enabling new capabilities such as instruction following, semantic scene understanding, and object-centric manipulation in real-world environments [25, 26]. By leveraging their ability to reason jointly over visual and linguistic modalities, LMMs offer a promising path toward more adaptable, scalable, and intuitive robotic systems that can operate with minimal task-specific training and additional data collection.
@@
Despite notable advances in high-level robotic task planning and embodied reasoning, the application of LMMs to robotic grasping remains relatively underexplored. Most existing work leveraging LMMs in robotics focuses on semantic instruction following, policy learning, or general manipulation planning, rather than on fine-grained grasp localization and selection. Among the few exceptions, PhyGrasp [27] introduces an LMM that integrates 3D point clouds and language to inform grasping decisions based on estimated physical properties such as material, fragility, and mass. However, PhyGrasp requires detailed 3D shape modeling and the availability of physical property annotations. It also depends on large-scale data collection and does not explicitly address task-driven, human-like grasp selection from single-view visual inputs. Another emerging line of work employs LMMs to direct grasping based on human verbal instructions, providing external context and object knowledge to guide the robot’s actions [28, 29, 30, 31]. While these approaches support task-specific grasping, they depend on human guidance at inference time and often require pretraining on large, specialized grasp datasets, limiting their scalability for fully autonomous deployment. Moreover, they focus on selecting which object to grasp, rather than identifying task-driven, human-like grasp locations on the object itself.
@@
In this letter, we explore the capability of LMMs to infer feasible, task-relevant grasps from single images. We introduce ORACLE-Grasp (O-Grasp), a novel grasping framework (Figure 1) that treats an LMM as a semantic oracle to guide grasp selection. Rather than relying on low-level image features or geometric heuristics, O-Grasp reformulates the grasping problem as an iterative, prompt-driven reasoning process, enabling human-like, task-aware grasp predictions even for novel and amorphous objects. Our approach requires no data collection, dataset curation, or pretraining, operating in a fully zero-shot manner by leveraging an existing, general-purpose LMM. Through a dual-prompt approach, O-Grasp extracts both high-level scene context and localized candidate regions, framing grasp selection as a structured, context-driven decision rather than free-form generation tasks.
@@
The prompts are designed to be general and autonomous, allowing O-Grasp to operate without human intervention at inference time, in contrast to prior approaches that require human guidance through language commands. Critically, the use of chain-of-thought [32] reasoning initiated by the first prompt encourages the model to produce task-driven, human-like grasps that align with intuitive manipulation strategies. Although O-Grasp does not explicitly optimize for grasp stability, experimental results demonstrate that the predicted grasps are physically feasible and lead to high robot pick up success rates. Additionally, the framework integrates an early stopping mechanism to improve efficiency and a depth-based refinement step to enhance grasp reliability when depth information is available.
II Method
@@
TABLE I: The two prompts used to provide context and locate a region to grasp
@@
While state-of-the-art LMM offer powerful semantic understanding and zero-shot object recognition, they fall short when applied to fine-grained part localization for robotic grasping. Outputs such as bounding boxes and masks often lack the spatial precision to isolate specific graspable parts, show inconsistency in the exact region they highlight across similar inputs, and do not convey the orientation required for proper gripper alignment. For instance, a handle might be detected with a bounding box that is loosely aligned, oversized, or arbitrarily rotated, making it unclear how the gripper should be positioned. Even segmentation masks, while richer in detail than boxes, often yield imprecise or fragmented regions that shift with minor input changes, offering no consistent contour or graspable axis. These limitations undermine the reliability of grasp execution, where even small errors in location or angle can lead to failure. To address this, this section introduces our method, which restructures grasp localization as an iterative decision process, enabling accurate and robust grasp prediction without explicit training.
II-A Problem Definition
We address the task of enabling a robotic arm to grasp objects placed within its workspace. The robotic arm is equipped with an RGB-D camera which can be mounted on the robot’s end-effector or somewhere in the workspace with the required calibration. We assume a single object is observed by the camera within the workspace. Although this assumption simplifies the setup, it could be relaxed in future work by incorporating context-aware reasoning based on the task or user-provided verbal instructions. The robot has no prior knowledge of the observed objects or the object to be grasped. Given an RGB image $\mathbf{I}_{i}$ and corresponding depth image $\mathbf{D}_{i}$ taken by the camera where the object is observed, the robot is required to pick up the object with a task-driven grasp. It is important to note that a standard RGB camera can be used in place of an RGB-D camera; however, this excludes the possibility of depth-based grasp refinement, which is discussed in a later section. Compensating for the lack of depth data would require additional techniques, such as visual servoing, which are beyond the scope of this work.
II-B Grasp Candidate Generation
LMMs, while proficient at generating natural language descriptions and answering image-related queries, exhibit limitations in spatial precision when embedding images into their latent space. This deficiency hinders their ability to produce pixel-accurate oriented bounding boxes for specific object parts, such as ’the handle of the mug’ or ’the handle of the drill’. Direct bounding box generation requests often yield imprecise or inconsistent outputs. Hence, we propose a novel strategy within the O-Grasp framework to guide the LMM. We first define the image tiling operation $\Gamma_{a\times b}(\mathbf{I}_{j})$ as the process of discretizing the image $\mathbf{I}_{j}$ into an $a\times b$ grid of equally sized tiles. In doing so, the problem becomes a multiple–choice question of candidate regions, where the LMM can select a tile instead of a free–form generation task. This approach is particularly effective for objects without obvious handles (e.g., a cardboard, a wallet or a shoe) and leads to more human–like grasp selections.
@@ II-B Grasp Candidate Generation
Figure 2: Illustration of a grasp candidate generation given an image $\mathbf{I}_{i}$ of a remote control object. Red cells on the grid denote selection for grasping by the LMM. After $m=3$ iterations with the GRP prompt over different grid configurations, stopping criterion (2) is satisfied leading to an additional $m$ iterations on a cropped image. Finally, the position $\bar{\mathbf{p}}$ and orientation angle $\theta$ of the required grasp is outputted.
@@
Dual-prompt masking. To guide the LMM, we propose the use of a dual-prompt process that leverages tool calling capabilities. The first prompt, referred to as the Scene Context Prompt (SCP), is paired with the original image $\mathbf{I}_{i}$ and instructs the model to generate a concise description of the primary object in the scene, with an emphasis on grasp-relevant features. Here, the SCP acts as a tool call that extracts high-level semantic context, helping the model to reason more effectively about where and how to grasp the object. An example of this prompt is shown in Table I.a. This step encourages the model to engage in a chain-of-thought reasoning process about the scene, producing a contextual response that subsequently informs the subsequent prompt. The second prompt, termed Grasp Region Prompt (GRP) and seen in Table I.b, aims to approximate a grasp location. To overcome the spatial imprecision of LMMs and facilitate effective reasoning over grasp locations, we discretize the image space into a set of discrete candidate areas. It is rolled out in multiple iterations to acquire a probabilistic estimation of the grasp region. In each iteration, the GRP is paired with an image $\Gamma_{u\times v}(\mathbf{I}_{i})$ such that $u,v\in\{3,4,\dots,9\}$. The values for $v$ and $u$ are selected without repetition, starting with coarse grids and gradually refined through additional GRP queries. The GRP incorporates the contextual output of the SCP to instruct the LMM to choose the optimal tile from the discretized image (Figure 2). Repeating the GRP procedure over $K$ iterations with varying grid configurations produces a set of candidate regions =$\mathcal{B}=\{B_{1},B_{2},\dots,B_{K}\}$, where $B_{j}$ denotes a rectangular pixel mask over the image, selected by the LMM during iteration $j$. Discretizing the image space into a grid of candidate areas enables a more robust and simple selection of grasp regions.
@@
Grasp position and orientation. For each pair of overlapping rectangular masks $B_{j}\in\mathcal{B}$ and $B_{k}\in\mathcal{B}$, we compute their Intersection-over-Union (IoU), denoted as $C_{ik}=IoU(B_{j},B_{k})$. If $C_{ik}>\gamma$, for a predefined threshold $\gamma$, the intersection region $B_{j}\cap B_{k}$ is added to the $\mathcal{B}$. A higher IoU value suggests a more promising grasp region, prompting the inclusion of an additional rectangle that reinforces areas with high grasp potential. After $K$ iterations and the inclusion of $s$ overlapping regions, the set $\mathcal{B}$ contains a total of $K+s$ rectangular pixel masks. In the subsequent step, the center coordinate $\mathbf{p}_{j}\in\mathbb{N}^{2}$ of each mask $B_{j}\in\mathcal{B}$ is computed and collected into a set $\mathcal{M}$, representing the centers of all masks. The centroid $\bar{\mathbf{p}}$ of set =$\mathcal{M}=\{\mathbf{p}_{1},\ldots,\mathbf{p}_{K+s}\}$ determines the final grasp point. Furthermore, Principal Component Analysis (PCA) is applied to the set of center points in $\mathcal{M}$ to extract the principal eigenvector $\mathbf{v}^{*}=(v^{*}_{x},v^{*}_{y})$ associated with the largest eigenvalue. The grasp orientation angle $\theta$ can be extracted from

### v2

# ORACLE-Grasp: Zero-Shot Affordance-Aligned Robotic Grasping using Large Multimodal Models

@@ Abstract
Grasping unknown objects in unstructured environments is a critical challenge for service robots, which must operate in dynamic, real-world settings such as homes, hospitals, and warehouses. Success in these environments requires both semantic understanding and spatial reasoning. Traditional methods often rely on dense training datasets or detailed geometric modeling, which demand extensive data collection and do not generalize well to novel objects or affordances. We present ORACLE-Grasp, a zero-shot framework that leverages Large Multimodal Models (LMMs) as semantic oracles to guide affordance-aligned grasp selection, without requiring task-specific training or manual input. The system reformulates grasp prediction as a structured, iterative decision process, using a dual-prompt tool-calling strategy: the first prompt extracts high-level object semantics, while the second identifies graspable regions aligned with the object’s function. To address the spatial limitations of LMMs, ORACLE-Grasp discretizes the image into candidate regions and reasons over them to produce human-like and context-sensitive grasp suggestions. A depth-based refinement step improves grasp reliability when available, and an early stopping mechanism enhances computational efficiency. We evaluate ORACLE-Grasp on a diverse set of RGB and RGB-D images featuring both everyday and AI-generated objects. The results show that our method produces physically feasible and semantically appropriate grasps that align closely with human annotations, achieving high success rates in real-world pick-up tasks. Our findings highlight the potential of LMMs for enabling flexible and generalizable grasping strategies in autonomous service robots, eliminating the need for object-specific models or extensive training.
keywords:
Grasping, Large Multimodal Models, Affordance-aligned Grasping
1 Introduction
Robotic grasping remains a core challenge in autonomous manipulation, particularly for service robots operating in unstructured and dynamic environments such as homes, hospitals or warehouses. Traditional approaches often depend on precise object models, depth-based geometry recognition, or supervised learning pipelines trained on large-scale datasets [Eppner2015, Levine2018]. However, real-world deployment frequently requires interacting with novel objects that exhibit unpredictable shapes and require task-based reasoning. In such scenarios, accurate grasp localization, identifying grasp points and orientations that are both physically feasible and task-appropriate, becomes critical [Zeng2018]. While recent deep learning methods have advanced grasp quality prediction [Mahler2019], they primarily focus on geometric properties and neglect semantic reasoning about the object or intended task. This gap motivates the development of new frameworks that combine high-level semantic understanding with spatial reasoning, enabling more human-like, adaptive grasp selection in the real-world.
@@ 1 Introduction
Previous approaches have typically relied on geometric analysis of depth data [Bohg2014, Huang2024], model-based grasp synthesis [Sahbani2012] or extensive training of deep networks on large-scale grasp datasets [Morrison2018, Khor2024]. Early work on learning-based methods for grasping novel objects used machine learning techniques to predict effective grasp points directly from 2D images without requiring complete 3D reconstruction or object recognition [Saxena2008]. However, deep learning-based approaches typically focus on learning and predicting the quality of candidate grasps, without integrating semantic reasoning about the task or the object’s context [Mahler2017]. While recent methods have made significant advances in pixel-wise grasp prediction and affordance learning [Shao2020, Zeng2022], they often lack the semantic reasoning needed for affordance-aligned grasping, especially in zero-shot settings where the object has not been seen before. Grasp prediction in real-world contexts often involves more than simply securing a stable hold; it requires selecting grasp points that are appropriate for the intended use of the object. For example, while a hammer can be grasped by either the head or the handle from a geometric perspective, semantically meaningful use (e.g., hammering a nail or handing it to a human) strongly favors the handle. Semantic reasoning helps resolve such ambiguities by considering object function and affordance, which are crucial for selecting grasps that are not only physically robust but also task-relevant and human-like [Dang2012]. Without this layer of understanding, grasping methods risk producing functionally incorrect or unsafe outcomes.
Large Language Models (LLMs) [Minaee2024] and Large Multimodal Models (LMMs) [wang2024] have revolutionized robotics by enabling advanced natural language understanding and multimodal perception [wang2024large]. These powerful models empower robots to interpret complex instructions, reason about their environment, and perform tasks with greater autonomy and flexibility. Despite notable advances of LMMs in high-level robotic task planning and embodied reasoning, their application to robotic grasping remains less explored. Most existing work leveraging LMMs in robotics focuses on semantic instruction following, policy learning or general manipulation planning, rather than on fine-grained grasp localization and selection.
@@
Existing approaches include PhyGrasp [guo2024], which introduces an LMM that integrates 3D point clouds and language to inform grasping decisions based on estimated physical properties such as material, fragility and mass. However, PhyGrasp requires detailed 3D shape modeling and the availability of physical property annotations. It also depends on large-scale data collection and does not explicitly address task-driven, human-like grasp selection from single-view visual inputs. Another emerging line of work employs LMMs to direct grasping based on human verbal instructions, providing external context and object knowledge to guide the robot’s actions [Tziafas2023, Jin2024, Xu2024, li2024, Li2024ShapeGrasp]. While these approaches support task-specific grasping, they depend on human guidance at inference time and often require pretraining on large, specialized grasp datasets, limiting their scalability for fully autonomous deployment. Moreover, they focus on selecting which object to grasp, rather than identifying affordance-aligned, human-like grasp locations on the object itself. Furthermore, these methods often rely on supervised components for grasp detection and quality evaluation, lack explicit spatial reasoning, and exhibit limited generalizability as they are typically trained on a fixed set of object classes, thereby hindering truly zero-shot execution [Tang2023b, Tang2025].
In this work, we investigate the ability of LMMs to generate context-free affordance-aligned grasps, i.e., grasps that do not rely on explicit object or task grounding provided by the user, but instead reflect the most probable or conventional way an object would be grasped to perform its likely function. For example, a knife or drill is typically grasped by the handle, regardless of explicit instruction. Unlike prior approaches that require structured input or custom-trained grasp encoders [minderer2022simple, Tang2023b, shridhar2023perceiver], we explore whether existing pretrained foundation models can reason about unfamiliar objects and infer meaningful grasps without collecting any task-specific training data or performing additional training. This direction enables more autonomous and flexible robot behavior, where high-level user commands can be interpreted without detailed object specifications, and the robot can independently identify and manipulate suitable objects. To focus on this capability, we limit our study to scenes containing a single unfamiliar object, allowing us to isolate the model’s ability to perform semantic and spatial reasoning. Additionally, we aim to assess how well LMMs can localize functional regions and support spatially grounded grasp decisions in the absence of dense supervision or structured priors.
@@
In this paper, we explore the capability of LMMs to infer feasible, task-relevant grasps from single images. We introduce ORACLE-Grasp, a novel grasping framework (Figure 1) that treats an LMM as a semantic oracle to guide grasp selection. Rather than relying on low-level image features or geometric heuristics, ORACLE-Grasp reformulates the grasping problem as an iterative, prompt-driven reasoning process, enabling human-like, task-aware grasp predictions even for novel and amorphous objects. Our approach requires no data collection, dataset curation or pretraining, operating in a fully zero-shot manner by leveraging an existing, general-purpose LMM. Through a dual-prompt approach, ORACLE-Grasp extracts both high-level scene context and localized candidate regions, framing grasp selection as a structured, context-driven decision rather than free-form generation tasks.
@@
The prompts are designed to be general and autonomous, allowing ORACLE-Grasp to operate without human intervention at inference time, in contrast to prior approaches that require human guidance through language commands. Critically, the use of chain-of-thought [Wei2022] reasoning initiated by the first prompt encourages the model to produce affordance-aligned, human-like grasps that align with intuitive manipulation strategies. Although ORACLE-Grasp does not explicitly optimize for grasp stability, experimental results demonstrate that the predicted grasps are physically feasible and lead to high robot pick-up success rates. Additionally, the framework integrates an early stopping mechanism to improve efficiency and a depth-based refinement step to enhance grasp reliability when depth information is available.
To summarize, our main contributions are as follows:
  • We introduce ORACLE-Grasp, a novel zero-shot framework for affordance-aligned grasping that leverages an LMM as a semantic oracle to infer human-like grasps without requiring training, grasp labels or custom encoders.
  • We propose a dual-prompting strategy that enables structured, chain-of-thought reasoning about object function and grasp region suitability directly from a single RGB image, without external task or object grounding.
  • We explore the ability of LMMs to perform context-free grasping, isolating their capacity to reason about unfamiliar, single-object scenes without requiring user-provided task annotations or structured priors.
  • We show that ORACLE-Grasp supports both semantic and spatial reasoning using only pretrained models, and demonstrate how depth data can be optionally integrated to refine and improve physical grasp execution.
  • We validate our approach in real-world robotic experiments across a variety of novel, untrained objects, demonstrating robust task-relevant grasping in fully zero-shot settings.
2 Related Work
2.1 Large Language and Multimodal Models
LLMs have significantly advanced robotics by enabling natural language understanding for task planning, human-robot interaction and high-level decision-making [ahn2022doasido, shridhar2023cliport]. These models allow robots to interpret abstract instructions (e.g., ”organize the desk”) and generate structured action sequences with explainable reasoning. Building on these capabilities, recent developments have introduced LMMs, which integrate vision and language processing to tackle embodied tasks [brohan2023rt2, driess2023palme]. By jointly reasoning over visual inputs and textual prompts, LMMs extend the functionality of LLMs, enabling robots to perceive, interpret and act within their environments with greater semantic understanding.
Prominent examples of LMMs include GPT-4o [openai2024gpt4o], Flamingo [Alayrac2022] and LLaVA [liu2023llava], which have demonstrated impressive capabilities in grounding visual scenes through natural language reasoning [wang2024]. These models are typically trained on large-scale datasets comprising paired image and text data, allowing them to align visual and linguistic representations within a shared semantic space. As a result, LMMs can generalize across domains, leverage contextual information, and infer high-level semantic relationships that are critical for flexible and task-driven robotic behaviors [Zhang2024]. Recently, LMMs have begun to be applied in robotics to bridge perception, reasoning and action [brohan2023rt2, long2024, MonWilliams2025], enabling new capabilities such as instruction following, semantic scene understanding and object-centric manipulation in real-world environments [singh2023progprompt, lykov2024]. By leveraging their ability to reason jointly over visual and linguistic modalities, LMMs offer a promising path toward more adaptable, scalable and intuitive robotic systems that can operate with minimal task-specific training and additional data collection.
2.2 Object and Task Grounding
In robotic systems, grounding refers to the process of associating abstract symbols, such as natural language commands, with perceptual representations and executable actions. Two critical forms of grounding are object grounding, linking linguistic or symbolic references to specific objects in the environment, and task grounding, connecting high-level task descriptions to robot actions or action plans. Early work in object grounding focused on mapping referring expressions to visual objects using either handcrafted features or statistical models [tellex2011understanding]. With the rise of deep learning and large-scale vision-language models, more recent methods have enabled open-vocabulary grounding using models like CLIP [Radford2021] and OWL-ViT [minderer2022simple], which align visual and linguistic representations in a shared embedding space. These models enable zero-shot recognition and grounding of previously unseen object categories based on semantic similarity.
Task grounding goes a step further by interpreting full instructions and mapping them to parameterized actions or policies. Frameworks such as SAYCAN [ahn2022doasido] and PaLM-E [driess2023palme] use LLMs to parse user intent and compose low-level skills or motion plans accordingly. These approaches demonstrate the ability to decompose tasks like ”clean the table” into structured action sequences and ground them in sensor inputs. More recent work explores grounding in embodied agents through multimodal interaction [shridhar2023perceiver] and active learning [singh2023progprompt], which enables context-aware task understanding and execution in dynamic environments.
Despite these advances, grounding in open-world settings remains challenging due to ambiguity in language, sensor noise and the combinatorial complexity of real-world object-task-action relationships. A key limitation lies in the lack of fine-grained spatial reasoning where many models can associate text with object categories but struggle to localize specific object parts or understand spatial configurations relevant to physical interaction (e.g., “grasp the handle” vs. “pour from the spout”). This impairs the ability of robots to ground language in actionable spatial representations, especially when operating with partial observations, occlusions, or novel geometries. Ongoing research continues to explore how these models can bridge this gap by integrating semantic reasoning with geometric understanding in a generalizable and interpretable manner.
2.3 Using LLM and LMM for Robot Grasping
Several recent approaches leverage LLMs and LMMs for robot grasping. GraspCLIP [Tang2023] enables task-oriented grasping by combining user verbal input (object and task) with visual input for grounding. However, it requires custom task-labeled datasets, limiting scalability and zero-shot generalization, and it is unclear whether it uses depth data for spatial reasoning. Unlike GraspCLIP, our proposed ORACLE-Grasp approach is domain-agnostic. Similarly, GraspGPT [Tang2023b] is a task-oriented grasping framework that uses LLMs to generalize to novel object-task pairs by generating semantic descriptions from natural language instructions. While effective, GraspGPT relies on language input and learned grasp priors at inference, making it less suitable for autonomous, vision-only, zero-shot grasping compared to ORACLE-Grasp. Additionally, GraspGPT evaluates grasp candidates using learned scoring, which might prioritize grasp success over task intent.
More recently, Grasp-Anything++ [Vuong2024] was presented as a large-scale language-driven grasp detection framework utilizing a diffusion model trained on over 10 million synthetic language-annotated grasp samples. It enables zero-shot performance guided by natural language instructions, however, its reliance on extensive training data and language input at inference time constrains its autonomy. Unlike ORACLE-Grasp, it lacks interpretable spatial reasoning and geometric refinement. LAN-grasp [Mirjalili2024] is a zero-shot semantic grasping method that integrates an LLM, LMM and a conventional grasp planner. The LLM identifies the appropriate object part from a user-provided label, which the LMM grounds visually, and a grasp planner generates constrained grasps. However, LAN-grasp’s dependence on external object labels and full 3D reconstruction limits its autonomy and real-time applicability.
ShapeGrasp is a zero-shot task-oriented grasping method that decomposes novel objects into simple geometric parts, represented as graphs [Li2024ShapeGrasp]. An LLM assigns semantics and selects the most suitable part to grasp based on the object name and task. The grasp is then executed on the part deemed most suitable, using basic heuristics over the part’s 3D geometry. However, ShapeGrasp relies heavily on handcrafted geometric decomposition, segmentation and symbolic reasoning, limiting its adaptability to complex objects. It lacks direct visual grounding and spatial precision, as it does not integrate LMM or learn from dense perceptual data. Unlike our approach, which performs end-to-end grounding and learns spatially-informed affordance-aligned grasps, ShapeGrasp depends on manual heuristics and coarse geometry, reducing its generality and robustness in cluttered or ambiguous scenes.
FoundationGrasp [Tang2025] is another task-oriented grasping framework that utilizes LLMs and VLMs for generalization to novel objects. It combines semantic and geometric reasoning through LLM descriptions, LLM/VLM-based encoders and a Transformer for grasp evaluation. However, FoundationGrasp requires task-specific user input and learned grasp evaluation networks, which restricts its autonomy and zero-shot capabilities. Unlike ORACLE-Grasp, it relies on supervised components and lacks interpretable, iterative spatial reasoning, potentially reducing its generalizability. Since LMMs typically lack precise spatial reasoning, these approaches often rely on custom-trained grasp detection models, which constrain their ability to generalize to new objects and tasks.
3 Methods
While state-of-the-art LMM offer powerful semantic understanding and zero-shot object recognition, they fall short when applied to fine-grained part localization for robotic grasping. Outputs such as bounding boxes and masks often lack the spatial precision to isolate specific graspable parts, show inconsistency in the exact region they highlight across similar inputs, and do not convey the orientation required for proper gripper alignment. For instance, a handle might be detected with a bounding box that is loosely aligned, oversized or arbitrarily rotated, making it unclear how the gripper should be positioned. Even segmentation masks, while richer in detail than boxes, often yield imprecise or fragmented regions that shift with minor input changes, offering no consistent contour or graspable axis. These limitations undermine the reliability of grasp execution, where even small errors in location or angle can lead to failure. To address this, this section introduces our method, which restructures grasp localization as an iterative decision process, enabling accurate and robust grasp prediction without explicit training.
3.1 Problem Definition
We address the task of affordance-aligned grasp prediction for unfamiliar objects using a robotic arm equipped with an RGB-D camera. The camera may be mounted on the robot’s end-effector or statically placed in the workspace, assuming proper calibration. We assume a simplified setting where a single, previously unseen object is present in the robot’s workspace. This setup allows us to isolate and study the semantic and spatial reasoning capabilities of LMMs. In future work, this assumption could be relaxed by incorporating context-aware reasoning based on the task or user-provided verbal instructions. It is important to note that a standard RGB camera can be used in place of an RGB-D camera; however, this excludes the possibility of depth-based grasp refinement, which is discussed in a later section. Compensating for the lack of depth data would require additional techniques, such as visual servoing, which are beyond the scope of this work.
The input to the system is an RGB image $\mathbf{I}_{i}$ and a corresponding depth map $\mathbf{D}_{i}$ of the scene, both captured from a single viewpoint. The object of interest is assumed to be visible in the image, without requiring prior segmentation or labeling. The system predicts a single affordance-aligned grasp pose for the object, defined as a grasp tuple $(\bar{\mathbf{p}}_{i},\theta_{i})$, where $\bar{\mathbf{p}}_{i}\in\mathbb{N}^{2}$ is the grasp position and $\theta_{i}\in\mathbb{R}$ is the required gripper orientation angle, both with respect to the image coordinate frame. This grasp should not only be physically feasible but also semantically meaningful, aligning with the object’s most likely functional usage (e.g., grasping a knife by the handle rather than the blade). The goal is to perform this prediction in a zero-shot setting: no object labels, segmentation masks, training data or prior knowledge about object geometry or category are provided.
3.2 Grasp Candidate Generation
LMMs, while proficient at generating natural language descriptions and answering image-related queries, exhibit limitations in spatial precision when embedding images into their latent space. This deficiency hinders their ability to produce pixel-accurate oriented bounding boxes for specific object parts, such as ’the handle of the mug’ or ’the handle of the drill’. Direct bounding box generation requests often yield imprecise or inconsistent outputs. Hence, we propose a novel strategy within the ORACLE-Grasp framework to guide the LMM. We first define the image tiling operation $\Gamma_{a\times b}(\mathbf{I}_{j})$ as the process of discretizing the image $\mathbf{I}_{j}$ into an $a\times b$ grid of equally sized tiles. In doing so, the problem becomes a multiple–choice question of candidate regions, where the LMM can select a tile instead of a free–form generation task. This approach is particularly effective for objects without obvious handles (e.g., a cardboard, a wallet or a shoe) and leads to more human–like grasp selections.
@@ 3.2 Grasp Candidate Generation
Table 1: The two prompts used to provide context and locate a region to grasp
@@
Figure 2: Illustration of a grasp candidate generation given an image $\mathbf{I}_{i}$ of a remote control object. Red cells on the grid denote selection for grasping by the LMM. After $m=3$ iterations with the GRP prompt over different grid configurations, stopping criterion (2) is satisfied leading to an additional $m$ iterations on a cropped image. Finally, the position $\bar{\mathbf{p}}$ and orientation angle $\theta$ of the required grasp are outputted.
@@
Dual-prompt masking. To guide the LMM, we propose the use of a dual-prompt process that leverages tool calling capabilities. The first prompt, referred to as the Scene Context Prompt (SCP), is paired with the original image $\mathbf{I}_{i}$ and instructs the model to generate a concise description of the primary object in the scene, with an emphasis on grasp-relevant features. Here, the SCP acts as a tool call that extracts high-level semantic context, helping the model to reason more effectively about where and how to grasp the object. An example of this prompt is shown in Table 1.a. This step encourages the model to perform Chain-of-Thought (CoT) reasoning about the scene, generating a semantic context that guides the following prompt’s spatial selection. The second prompt, termed Grasp Region Prompt (GRP) and seen in Table 1.b, aims to approximate a grasp location. To overcome the spatial imprecision of LMMs and facilitate effective reasoning over grasp locations, we discretize the image space into a set of discrete candidate areas. It is rolled out in multiple iterations to acquire a probabilistic estimation of the grasp region. In each iteration, the GRP is paired with an image $\Gamma_{u\times v}(\mathbf{I}_{i})$ such that $u,v\in\{3,4,\dots,9\}$. The values for $v$ and $u$ are selected without repetition, starting with coarse grids and gradually refined through additional GRP queries.
The GRP incorporates the contextual output of the SCP (i.e., <SCP response>) to instruct the LMM to choose the optimal tile from the discretized image (Figure 2). Crucially, the GRP output includes a required <brief explanation of your choice> field, which serves as a mechanism for CoT reasoning (Table 1.b). By forcing the LMM to articulate its semantic and functional rationale before committing to a specific grid cell, the framework ensures the model actively applies contextual constraints to the spatial domain, such as prioritizing handles or avoiding fragile parts. This requirement acts as a vital bridge between the high-level affordances identified by the SCP and the low-level spatial grid, grounding the final grasp selection in a semantically-informed reasoning process rather than a purely geometric estimation. Furthermore, these explanations provide interpretability, allowing the zero-shot decision process to be verified against the object’s perceived functional properties. Repeating the GRP procedure over $K$ iterations with varying grid configurations produces a set of candidate regions $\mathcal{B}=\{B_{1},B_{2},\dots,B_{K}\}$, where $B_{j}$ denotes a rectangular pixel mask over the image, selected by the LMM during iteration $j$. Discretizing the image space into a grid of candidate areas enables a more robust and simple selection of grasp regions.
@@
Grasp position and orientation. For each pair of overlapping rectangular masks $B_{j}\in\mathcal{B}$ and $B_{k}\in\mathcal{B}$, we compute their Intersection-over-Union (IoU), denoted as $C_{ik}=IoU(B_{j},B_{k})$. If $C_{ik}>\gamma$, for a predefined threshold $\gamma$, the intersection region $B_{j}\cap B_{k}$ is added to the $\mathcal{B}$. A higher IoU value suggests a more promising grasp region, prompting the inclusion of an additional rectangle that reinforces areas with high grasp potential. After $K$ iterations and the inclusion of $s$ overlapping regions, the set $\mathcal{B}$ contains a total of $K+s$ rectangular pixel masks. In the subsequent step, the center coordinate $\mathbf{p}_{j}\in\mathbb{N}^{2}$ of each mask $B_{j}\in\mathcal{B}$ is computed and collected into a set $\mathcal{M}$, representing the centers of all masks. The centroid $\bar{\mathbf{p}}$ of set $\mathcal{M}=\{\mathbf{p}_{1},\ldots,\mathbf{p}_{K+s}\}$ determines the final grasp point. Furthermore, Principal Component Analysis (PCA) is applied to the set of center points in $\mathcal{M}$ to extract the principal eigenvector $\mathbf{v}^{*}=(v^{*}_{x},v^{*}_{y})$ associated with the largest eigenvalue. The grasp orientation angle $\theta$ can be extracted from